# Unified Multimodal Decoding of Brain Signals

### v1

@@ Abstract
We address prevailing challenges of the brain-powered research, departing from the observation that the literature hardly recover accurate spatial information and require subject-specific models. To address these challenges, we propose UMBRAE, a unified multimodal decoding of brain signals. First, to extract instance-level conceptual and spatial details from neural signals, we introduce an efficient universal brain encoder for multimodal-brain alignment and recover object descriptions at multiple levels of granularity from subsequent multimodal large language model (MLLM). Second, we introduce a cross-subject training strategy mapping subject-specific features to a common feature space. This allows a model to be trained on multiple subjects without extra resources, even yielding superior results compared to subject-specific models. Further, we demonstrate this supports weakly-supervised adaptation to new subjects, with only a fraction of the total training data. Experiments demonstrate that UMBRAE not only achieves superior results in the newly introduced tasks but also outperforms methods in well established tasks. To assess our method, we construct and share with the community a comprehensive brain understanding benchmark BrainHub. Our code and benchmark are available at https://github.com/weihaox/UMBRAE.
@@ 1 Introduction
However, there are remaining challenges in brain-powered research. First, decoding brain signals into a single modality results in a lossy compression of the brain activities. On the one hand, text fails to preserve the peculiar appearance of a texture or the spatial location of an object. On the other hand, visual decoding [53, 34, 41] addresses the underdetermined problem of pixel-wise reconstruction and lacks explicitation of the scene structure. Consider, for instance, the scenario where a person uses thoughts to control a robotic arm to retrieve an apple from a fruit bowl on a table. The first task is to recognize the apple amidst similar visual concepts and then locate its exact position. But current methods lack such fine-grained decoding capability to interpret object categories, visual concepts, and their relationships. The second challenge pertains to the subject-specific patterns of brain activities [2]. Therefore, current methods typically train a model for each subject to cope with distinctive brain patterns. Decoding brain signals across multiple subjects presents challenges due to the structural and functional differences among individual brains.
@@
  • We introduce UMBRAE, aiming at unified multimodal decoding of brain responses. Our method relies on a universal brain encoder and a frozen multimodal large language model seeking to align brain signals with images. We also propose a cross-subject training strategy to learn a universal representation across subjects, as opposed to the standard subject-specific training. Furthermore, it allows the novel weakly-supervised adaptation, enabling the training of a model for any arbitrary subject with minimal training data.
@@ 3 UMBRAE
The overview of our method, named UMBRAE, is in Fig. 2. The acronym stands for unified multimodal decoding of brain responses and signifies the process of unveiling encoded information hidden within the ‘shadows’ of brain signals. We rely mainly on a flexible brain encoder architecture (Sec. 3.1) and a cross-subject training strategy (Sec. 3.2) to map brain responses from different subjects, each with variable length, into a common feature space. In our experiments, we observe that simply binding brain modality with image features enables the recovery of both semantic and spatial cues (Sec. 3.3). We then conduct brain prompting interface by inputting the standardized brain representations from diverse subjects into MLLMs for downstream tasks (Sec. 3.4).
@@ 3.1 Architecture
Our brain encoder is a flexible model architecture based on a lightweight transformer [50, 20] that accepts variable-length inputs of brain responses. This is important for cross-subject training as fMRI data encodes signals primarily from the most active brain regions, which are variable across subjects. Hence, our architecture comprises subject-specific tokenizers which aim is to extract and map the subject-specific characteristics, along with a universal perceive encoder designed to capture subject-agnostic knowledge, later aligned with image features.
@@ 3.3 Multimodal Alignment
Previous multimodal methods have learned to map multiple modalities into a common latent space, typically through contrastive pretraining, using either images [15] or text [16] as the binding modality. To preserve spatial information and improve generalizability, we align our brain representations with image features from a pretrained image encoder using element-wise reconstruction.
@@
Our brain encoder $\mathcal{B}$ learns the alignment between source brain space $\mathbf{B}$ and target image space $\mathbf{V}$. Different from previous methods, we align the brain signals with the intermediate image features from a pretrained image encoder, thus achieving semantic and spatial alignment for the brain representation.
@@ 3.4 Brain Prompting Interface
After alignment, brain features predicted by the brain encoder $\mathcal{B}$ are forwarded to the MLLM’s adapter, to retrieve the mapped visual embeddings <image>. These embeddings are then concatenated with a user instruction prompt and inputted into the finetuned LLM. Thus, our brain encoder inherits from multimodal capabilities of the MLLM, allowing tasks to be used in a prompting fashion using template:
@@
Our primary focus in this study is on brain captioning and grounding, which reflects the capabilities of brain signals in concept recognition and spatial localization. These two tasks are often referred to as image captioning and visual grounding in the MLLM literature [30, 9, 48]. However, in this context, brain signals are used as input rather than images. Our method also supports other instruction-following capability, such as conversation, detail description, and complex reasoning. Our method is model-agnostic, allowing for the use of any image encoders, large language models, and multimodal large language models according to specific needs.
@@ 4.2 BrainHub
Table 1: Brain Captioning. ‘UMBRAE-S1’ refers to our model trained with a single subject (S1 here) only, while ‘UMBRAE’ denotes the model with cross-subject training. ‘Shikra-w/img’ refers to the image captioning result from Shikra [9] using the ground truth image as input, serving as an approximate upper bound. The colors represent the best, second-best, and third-best performance. Results for the other subjects are provided in the appendix.
@@ 5.2 Training Strategies
Current vision-language models typically comprise three main components: an image encoder, an adapter, and a large language model. Within this framework, there are several potential ways for multimodal-brain alignment. For example, one could train the model to map brain responses to a pretrained semantic space through contrastive learning, which has been a common practice in previous methods. Alternatively, one could opt to fine-tune the adapter or train it jointly with the brain encoder, applying separate losses to each component. In this section, we delve into the motivation and rationale behind aligning with image features from the image encoder using a reconstruction loss.
@@ Appendix 0.A UMBRAE: Tasks and Example Prompts
It should be noted that our method is compatible with nearly all tasks featured in Shikra [9] and LLaVA [30], with the exception of tasks that necessitate initially locating an object or scene as inputs, such as grounding captioning and referring expression generation. Importantly, note that although the term ‘image’ is used in the prompts we do not utilize images as input in our brain understanding tasks. The reference images are only used for visualization purposes. In practice, the prompt embedding will be concatenated with the brain features predicted from our brain encoder, utilizing brain signals as inputs.
We report detailed results for each task of Tab. 1 later in Appendix 0.C.
@@ 0.B.2 Benchmark Construction
Category
COCO classes (# of classes)
A
S + I (80)
S
SC + SO (28)
SC
person, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe (11)
SO
bicycle, car, motorcycle, airplane, bus, train, truck, boat, bench, chair, couch, bed, dining table, toilet, sink, refrigerator, clock (17)
I
traffic light, fire hydrant, stop sign, parking meter, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, potted plant, tv, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, book, vase, scissors, teddy bear, hair drier, toothbrush (52)
@@
Figure 1: BrainHub Statistics. We illustrate the statistics in (0(a)) and the mapping relationships in (0(b)) for the categories ontology used in BrainHub, w.r.t. to the original COCO classes. Please zoom in (0(a)) for details.
@@ 0.C.1 Brain Captioning
UMBRAE 56.94 37.83 24.99 16.92 17.99 41.84 52.26 11.51 64.66 70.98
Image (for reference only)
@@
SDRecon [45] some of the amazingly colorful african animals, the only because its natural habitat is
BrainCap [14] a close up of a person with a ball
OneLLM [16] A black and white sheep standing next to a green bush.
UMBRAE-S1 A small bird is sitting on a tree branch.
UMBRAE A bird is sitting on the branches of a tree.
@@
Image (for reference only)
@@
COCO [29] A man riding a snowboard down a hill.
Shikra-w/img [9] A skier is going down a snowy hill.
S1 A person in a ski outfit skiing down a slope.
S2 A man riding a surfboard on top of a wave.
S5 A person on skis is skiing on a snowy slope.
S7 A person riding a snowboard on top of a snow covered slope.
@@ 0.C.4 Visual Decoding
Fig. 8 displays a qualitative comparison of individual subjects, while the quantitative evaluation of UMBRAE on all four subjects can be found in Tab. 6.
Reference
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SD [40]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SDXL [37]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kandinsky3 [4]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GLIGEN [26]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD [55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-2
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-3
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Table 6: Subject-Specific Visual Decoding Evaluation. Quantitative evaluation of the UMBRAE reconstruction for the four subjects (S1, S2, S05, and S7) of NSD [2].
@@
Reference
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S2
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S5
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S7
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix 0.D Additional Analyses
Here, we provide a couple of additional studies to further evaluate the per-subject performance (Sec. 0.D.1), the weakly-supervised adaptation (Sec. 0.D.2). We also build on MLLMs capacities to explore other tasks (Sec. 0.D.3) and finally, demonstrate the model-agnostic characteristics of our method (Sec. 0.D.4).
@@ 0.D.2 Weakly-Supervised Adaptation
Figure 9: Weakly-Supervised Subject Adaptation. We plot our performance for adaptation to a given subject (S1 in 8(a), S2 in 8(b), S5 in 8(c)). For adapting to a given subject Sx, we first train UMBRAE on all other subjects and then finetune the pretrained model with varying ratios of training data from the target subject.
0.D.3 Other Supported Tasks

### v2

# UMBRAE: Unified Multimodal Brain Decoding

@@ Abstract
We address prevailing challenges of the brain-powered research, departing from the observation that the literature hardly recover accurate spatial information and require subject-specific models. To address these challenges, we propose UMBRAE, a unified multimodal decoding of brain signals. First, to extract instance-level conceptual and spatial details from neural signals, we introduce an efficient universal brain encoder for multimodal-brain alignment and recover object descriptions at multiple levels of granularity from subsequent multimodal large language model (MLLM). Second, we introduce a cross-subject training strategy mapping subject-specific features to a common feature space. This allows a model to be trained on multiple subjects without extra resources, even yielding superior results compared to subject-specific models. Further, we demonstrate this supports weakly-supervised adaptation to new subjects, with only a fraction of the total training data. Experiments demonstrate that UMBRAE not only achieves superior results in the newly introduced tasks but also outperforms methods in well established tasks. To assess our method, we construct and share with the community a comprehensive brain understanding benchmark BrainHub. Our code and benchmark are available at https://weihaox.github.io/UMBRAE.
@@ 1 Introduction
However, there are remaining challenges in brain-powered research. First, decoding brain signals into a single modality results in a lossy representation of the brain activities. On the one hand, text fails to preserve the peculiar appearance of a texture or the spatial location of an object. On the other hand, visual decoding [53, 34, 41] addresses the underdetermined problem of pixel-wise reconstruction and lacks explicitation of the scene structure. Consider, for instance, the scenario where a person uses thoughts to control a robotic arm to retrieve an apple from a fruit bowl on a table. The first task is to recognize the apple amidst similar visual concepts and then locate its exact position. But current methods lack such fine-grained decoding capability to interpret object categories, visual concepts, and their relationships. The second challenge pertains to the subject-specific patterns of brain activities [2]. Therefore, current methods typically train a model for each subject to cope with distinctive brain patterns. Decoding brain signals across multiple subjects presents challenges due to the structural and functional differences among individual brains.
@@
  • We introduce UMBRAE, aiming at unified multimodal brain decoding. Our method relies on a universal brain encoder and a frozen multimodal large language model seeking to align brain signals with images. We also propose a cross-subject training strategy to learn a universal representation across subjects, as opposed to the standard subject-specific training. Furthermore, it allows the novel weakly-supervised adaptation, enabling the training of a model for any arbitrary subject with minimal training data.
@@ 3 UMBRAE
The overview of our method, named UMBRAE, is in Fig. 2. The acronym stands for unified multimodal brain decoding and signifies the process of unveiling encoded information hidden within the ‘shadows’ of brain signals. We rely mainly on a flexible brain encoder architecture (Sec. 3.1) and a cross-subject training strategy (Sec. 3.2) to map brain responses from different subjects, each with variable length, into a common feature space. In our experiments, we observe that simply binding brain modality with image features enables the recovery of both semantic and spatial cues (Sec. 3.3). We then conduct brain prompting interface by inputting the standardized brain representations from diverse subjects into MLLMs for downstream tasks (Sec. 3.4).
@@ 3.1 Architecture
Our brain encoder, based on a lightweight transformer architecture [50, 20], accommodates variable-length brain response inputs. This is important for cross-subject training as fMRI data are variable across subjects. Hence, our architecture comprises subject-specific tokenizers which aim is to extract and map the subject-specific characteristics, along with a universal perceive encoder designed to capture subject-agnostic knowledge, later aligned with image features.
@@ 3.3 Multimodal Alignment
Previous multimodal methods learn to map multiple modalities into a common latent space, typically through contrastive pretraining, using either images [15] or text [16] as the binding modality. Instead, we align brain representations with image features from a pretrained image encoder using element-wise reconstruction.
@@
Our brain encoder $\mathcal{B}$ learns the alignment between source brain space $\mathbf{B}$ and target image space $\mathbf{V}$. Different from previous methods, we align the brain signals with the intermediate image features from a pretrained image encoder, thus achieving semantic and spatial alignment for the brain representation. MLLMs [9, 30] show such features provide sufficient visual cues for finetuning LMMs. Furthermore, aligning with intermediate image features allows direct input of aligned brain representations into the MLLM.
@@ 3.4 Brain Prompting Interface
After alignment, brain features from the brain encoder $\mathcal{B}$ are fed into the MLLM’s adapter, to retrieve the mapped visual embeddings <image>. These embeddings are then concatenated with a user instruction prompt and inputted into the finetuned LLM. Thus, our brain encoder inherits from multimodal capabilities of the MLLM, allowing tasks to be used in a prompting fashion using template:
@@
Our primary focus in this study is on brain captioning and grounding, which reflects the capabilities of brain signals in concept recognition and spatial localization. They are often referred to as image captioning and visual grounding in the multimodal learning literature. However, in this context, brain signals are used as the input rather than images. Our method also supports other instruction-following capabilities, such as conversation, detail description, and complex reasoning. Our method is model-agnostic, allowing for the use of any image encoders, LLMs, and MLLMs according to specific needs.
@@ 4.2 BrainHub
Table 1: Brain Captioning. ‘UMBRAE-S1’ refers to our model trained with a single subject (S1 here) only, while ‘UMBRAE’ denotes the model with cross-subject training. ‘Shikra-w/img’ refers to the image captioning result from Shikra [9] using the ground truth image as input, serving as an approximate upper bound. The colors represent the best, second-best, and third-best performance. Brain captioning results for the other subjects are provided in the appendix.
@@ 5.2 Training Strategies
Current vision-language models typically comprise three main components: an image encoder, an adapter, and a large language model. Within this framework, there are several potential ways for multimodal-brain alignment. For example, one could train the model to map brain responses to a pretrained semantic space through contrastive learning, which has been a common practice in previous methods. Alternatively, one could opt to fine-tune the adapter or train it jointly with the brain encoder, applying separate losses to each component. In this section, we delve into the motivation and rationale behind aligning with image features from the image encoder using a reconstruction loss. Further, in the appendix we ablate our sampling strategy.
@@ Appendix 0.A UMBRAE: Tasks and Example Prompts
It should be noted that our method is compatible with nearly all tasks featured in Shikra [9] and LLaVA [30], with the exception of tasks that necessitate initially locating an object or scene as inputs, such as grounding captioning and referring expression generation. Importantly, note that although the term ‘image’ is used in the prompts we do not utilize images as input in our brain understanding tasks. The reference images are only used for visualization purposes. In practice, the prompt embedding will be concatenated with the brain features predicted from our brain encoder, utilizing brain signals as inputs. We report detailed results for each task of Tab. 1 later in Appendix 0.C.
@@ 0.B.2 Benchmark Construction
Category COCO classes (# of classes)
A S + I (80)
S SC + SO (28)
SC person, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe (11)
SO bicycle, car, motorcycle, airplane, bus, train, truck, boat, bench, chair, couch, bed, dining table, toilet, sink, refrigerator, clock (17)
I traffic light, fire hydrant, stop sign, parking meter, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, potted plant, tv, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, book, vase, scissors, teddy bear, hair drier, toothbrush (52)
@@ A S + I (80)
Figure 2: BrainHub Statistics. We illustrate the statistics in (2(a)) and the mapping relationships in (2(b)) for the categories ontology used in BrainHub, w.r.t. to the original COCO classes. Please zoom in (2(a)) for details.
@@ 0.C.1 Brain Captioning
UMBRAE 57.20 38.30 25.49 17.13 18.29 42.16 52.73 11.63 65.90 71.83
Image
(for reference only)
@@
SDRecon [45] some of the amazingly colorful african animals, the only because its natural habitat is BrainCap [14] a close up of a person with a ball OneLLM [16] A black and white sheep standing next to a green bush. UMBRAE-S1 A small bird is sitting on a tree branch. UMBRAE A bird is sitting on the branches of a tree.
@@
Image
(for reference only)
@@
COCO [29] A man riding a snowboard down a hill. Shikra-w/img [9] A skier is going down a snowy hill. S1 A person in a ski outfit skiing down a slope. S2 A man riding a surfboard on top of a wave. S5 A person on skis is skiing on a snowy slope. S7 A person riding a snowboard on top of a snow covered slope.
@@ 0.C.4 Visual Decoding
Fig. 10 displays a qualitative comparison of individual subjects, while the quantitative evaluation of UMBRAE on all four subjects can be found in Tab. 6. Fig. 8 shows comparison on visual decoding between our method and the literature.
Figure 8: Visual Decoding Comparison between UMBRAE and the literature on NSD. All reconstructed images are from S1.
Reference Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SD [40] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SDXL [37] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kandinsky3 [4] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GLIGEN [26] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD [55] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
VD-3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Table 6: Subject-Specific Visual Decoding Evaluation. Quantitative evaluation of the UMBRAE reconstruction for the four subjects (S1, S2, S5, and S7) of NSD [2].
@@
Reference Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S5 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S7 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix 0.D Additional Analyses
Here, we provide a couple of additional studies to further evaluate the per-subject performance (Sec. 0.D.1), the weakly-supervised adaptation (Sec. 0.D.2). We analyze different sampling strategies in cross-subject training (Sec. 0.D.3) and present a joint grounding-decoding evaluation (Sec. 0.D.4). We also build on MLLMs capacities to explore other tasks (Sec. 0.D.5) and finally, demonstrate the model-agnostic characteristics of our method (Sec. 0.D.6).
@@ 0.D.2 Weakly-Supervised Adaptation
Figure 12: Weakly-Supervised Subject Adaptation. We plot our performance for adaptation to a given subject (S1 in 12(a), S2 in 12(b), S5 in 12(c)). For adapting to a given subject Sx, we first train UMBRAE on all other subjects and then finetune the pretrained model with varying ratios of training data from the target subject.
0.D.3 Sampling Strategies in Cross-Subject Training
There are different sampling strategies of subjects and data samples in the cross-subject training. Our batch of $B$ samples is made of $\theta\!\times{}\!B$ samples from the same user chosen according to users frequencies, while the remaining samples are then sampled from other users. This is illustrated on the left of Tab. 8 using, for simplicity, four users, $\theta\!=\!0.44$ and $B\!=\!16$. ‘Random’ means all subjects are randomly sampled, while ‘Stratified’ ensures that data samples from the four subjects are equal in number within a batch. ‘Ours-R’ and ‘Ours’ are the same for the dominant subject but differ in the sampling strategies for the remaining three subjects. Using a dominant subject per batch helps the model to learn intra-subject variations while being exposed to other subjects patterns to enhance inter-subject discrimination and alleviate catastrophic forgetting. Tab. 8 reports average metrics across users for ‘Random’, ‘Stratified’, ‘Ours-R’, and ‘Ours’ (using $\theta\!=\!0.50$, $B\!=\!256$). Ours outperforms all other sampling strategies.
Sampling Captioning Grounding
BLEU1 CIDEr METER RefCLIPS acc@0.5 (A) IoU (A) acc@0.5 (S) IoU (S)
Random 51.84 35.37 15.52 67.26 13.49 17.39 21.56 25.39
Stratified 58.91 55.83 18.94 72.69 17.31 20.34 27.08 29.03
Ours-R 58.02 55.02 18.68 72.09 15.22 18.84 23.58 26.74
Ours 59.09 57.76 19.24 72.96 18.03 20.63 28.22 29.32
Table 8: Comparison of Sampling Strategies. ‘Random’ means all subjects are randomly sampled, while ‘Stratified’ ensures that data from the four subjects are equal in number within a batch. ‘Ours-R’ and ‘Ours’ are the same when sampling from the selected dominant subject but differ for the remaining three subjects.
0.D.4 Joint Grounding-Decoding Evaluation
We visualize grounding results on the reference images (ground truth) to better assess their performance. Fig. 13 further shows grounding and reconstruction simultaneously to highlight their synergy. Results demonstrate that the two tasks are correlated. In some cases, although grounding is correct, reconstruction is inaccurate (e.g. surfer, giraffe, and skier are well located but misorriented). The first row presents reference images. The second row displays reconstructed images, which are generated using the decoded texts and groundings from the third row as inputs. The third row illustrates the spotting captioning results, where the coordinates for each mentioned object are omitted and instead visualized in color within the corresponding generated images shown in the second row.
Figure 13: Joint Visualization of Grounding and Reconstruction. The first row presents reference images. The second row displays reconstructed images generated using decoded texts and groundings from the third row. The coordinates in the third row for each mentioned object are omitted and visualized in color in the second row. Reconstruction may be inaccurate even with correct bounding boxes.
0.D.5 Other Supported Tasks